A news website reported (Feb. 17, 2020) that ridership of Karachi Metro Bus declined after a fare increase: “There were nearly four million fewer riders in December 2019, the first full month after the price of a token increased Rs. 25 to Rs.150, than in the previous December, a 4.3 percent decline.”
i. Use these data to estimate the price elasticity of demand for Karachi Metro Bus. (2 Marks)
ii. According to your estimate, what happens to the Karachi Metro Bus’s revenue when the fare rises? (0.5 Marks)
iii. Why might your estimate of the elasticity be unreliable?
1
Expert's answer
2020-11-30T16:37:56-0500

i. Elasticity of demand = percentage change in demand/percentage change in price

Percentage change in demand = 4.3 %

Percentage change in price:

25 – x

(150 – 25) – 100 %

"x = \frac{25 \times 100}{125} = 20" %

Elasticity of demand "= \frac{-4.3}{20} = -0.21"

Negative sign is because a rise in price caused demand to fall. Price and demand move in opposite directions.

ii. The demand is inelastic. Karachi Metro Bus’s revenue will increase.